OpenStudy (anonymous):

calc help, intergral (2-x)/sqrt(4-x^(2)) help step by step please

OpenStudy (ash2326):

We have \[\int \frac{2-x}{\sqrt{4-x^2}}dx\] \[\int \frac{2}{\sqrt{4-x^2}}-\frac{x}{\sqrt{4-x^2}}dx\] Do you get this step?

OpenStudy (anonymous):

yeah

OpenStudy (ash2326):

Do you know the derivative of \(\sin^{-1}x\)?

OpenStudy (anonymous):

thats 1/sqrt(1-x^(2))

OpenStudy (anonymous):

i believe

OpenStudy (anonymous):

the first set i recieved 2arcsin(x/2) but i dont remember how to do the one with x as the numerator

OpenStudy (ash2326):

good:) second is easy

OpenStudy (ash2326):

\[\int-\frac{x}{\sqrt{4-x^2}}dx\] put \[4-x^2=t\] so \[-2x dx=dt\] or \[\frac{dt}{-2x}=dx\] now our integral \[\int-\frac{x}{\sqrt{t}}\frac{dt}{-2x}\] we get \[\int \frac{dt}{2\sqrt t}dt\] I think you can do this, can' t you?

OpenStudy (anonymous):

yup then i got 1/2 ln | u or 4-x^(2)| + c

OpenStudy (anonymous):

i just add that with the first one and im fine correct.

OpenStudy (ash2326):

It's not 1/t, you won't get ln t it's \[\int \frac{1}{2 \underline {\sqrt t}}dt\]

OpenStudy (anonymous):

oh, ok i see that but now im stuck lol

OpenStudy (anonymous):

it wouldnt be 1/2arcsin(x/2) would it?

OpenStudy (ash2326):

The first one is almost correct it'd be \[\arcsin \frac x 2\]

OpenStudy (anonymous):

how is the first equation not 2arcsin(x/2)?

OpenStudy (ash2326):

Sorry you're correct with the first one it'd be \[2\arcsin \frac x2\]

OpenStudy (anonymous):

o, ok and the second one would be 1/2arcsin(x/2)?

OpenStudy (ash2326):

use the following for second \[\int x^n dx=\frac{x^{n+1}}{n+1}\] here we have \[\int \frac{1}{\sqrt t} dt\] here n=\(-\frac 12\)

OpenStudy (ash2326):

nope, please read the post for the second integral again

OpenStudy (anonymous):

ohhh im had a dumb moment, i can bring the 1/sqrt(t) up and make it t^(-1)dt that becomes t^(1/2) / (1/2) or 2t(1/2)

OpenStudy (ash2326):

if you bring it up, it'll be \(\large t^{-\frac 1 2}\)

OpenStudy (anonymous):

^

OpenStudy (ash2326):

oops, after integration you're right then \[2\sqrt t\]

OpenStudy (ash2326):

you got it?

OpenStudy (anonymous):

yeah my final looks like 2arcsin(x/2) - sqrt(u) or sqrt (4-x^(2)) + c

OpenStudy (ash2326):

good work:D Sorry to confuse you in between:(

OpenStudy (anonymous):

lol its alright i appreciate the help.

OpenStudy (ash2326):

thanks:)